# Demonstrators for Industrial Cyber-Physical System Research: A Requirements Hierarchy Driven by Software-Intensive Design

### v1

@@ Abstract.
One of the challenges apparent in the organisation of research projects is the uncertainties around the subject of demonstrators. A precise and detailed elicitation of the coverage for project demonstrators is often an afterthought and not sufficiently detailed during proposal writing. This practice leads to continuous confusion and a mismatch between targeted and achievable demonstration of results, hindering progress. The reliance on the TRL scale as a loose descriptor does not help either. We propose a demonstrator requirements elaboration framework aiming to evaluate the feasibility of targeted demonstrations, making realistic adjustments, and assist in describing requirements. In doing so, we define 5 hierarchical levels of demonstration, clearly connected to expectations, e.g., work package interaction, and also connected to the project’s industrial use-cases. The considered application scope in this paper is the domain of software-intensive systems and industrial cyber-physical systems. A complete validation is not accessible, as it would require application of our framework at the start of a project and observing the results at the end, taking 4–5 years. Nonetheless, we have applied it to two research projects from our portfolio, one at the early and another at the final stages, revealing its effectiveness.
@@ 1. Introduction
Research projects often include “demonstrators” as deliverables. It is often assumed that a demonstrator is inherently known to project members, with no structure or definitive description provided as part of the project proposal. Another valid reason for the absence of such a description is the initial level of uncertainty specific to research projects. After all, a research project is supposed to be high-level by definition. Questions such as: “What is there to be demonstrated?”, “How much of it can be demonstrated?”, “Given the a priori knowledge limitations resulting from unknowns in a research project, which level of demonstration is feasible?”, are familiar to most researchers. Most projects try to solve this by relying on the Technology Readiness Levels (TRL), which proves to be inadequate as it does not account for a demonstrator’s coverage. We discuss this in Section 3.4.
@@
Scope
The role of software is ever-increasing in industrial Cyber-Physical Systems (CPS), to an extent that we can consider modern instances as Software-Intensive Systems (SIS) (Hölzl:2008:ESIA, 1). This designation is backed by the fact that analytic solutions intended for industrial CPS, especially the likes of solutions providing capabilities such as Fault Detection and Identification, anomaly detection/identification (Ibidunmoye:2015:PADB, 2), and predictive maintenance, are fully based on software algorithms and data. In this paper, we focus on research project in the context of SIS. Further evaluations with more diverse projects could lead to adjustments, making our method suitable for a broader scope.
@@ 2. Background
2.1. TRL levels
The Technology Readiness Level (TRL) scale in its original form, was first formalised by NASA (Mankins:2004:TRL, 3). The scale is intended as a framework for systematic assessment of technology maturity during its development. Consisting of 9 levels from basic research (TRL 1) to fully operational (TRL 9), it was primarily focused on hardware for aerospace technologies. These levels are defined as (taken from (Mankins:2004:TRL, 3)):
@@ 2.3.1. Stakeholders and building blocks
  • Industrial partners usually provide real-world use cases, validation platforms, or deployment opportunities.
@@
Demonstrator-related misalignments typically map to misunderstood, underspecified, or unspecified requirements. Especially in software-intensive projects, the absence of clearly defined interfaces, data formats, or operational contexts between WPs leads to friction during integration phases. Therefore, Requirement Engineering (RE) practices in research contexts should explicitly address targeted demonstrator scope and maturity.
@@ 2.3.2. WP dependencies
Considering the four major dependencies between software components, i.e., data dependency, control dependency, functional dependency, and temporal dependency, the most relevant dependencies for WPs can be considered as data and temporal. While we cannot dismiss the rest, especially given the possibility of per WP software converging into a unified platform, data and temporal dependencies stand out. Incorporation of control or functional dependencies is less common as these do not align with having separate WPs. Most WPs produce output for or consume input from other WPs, which also dictates order. Order of processing can be present without any exchange as well, e.g., an analysis requiring an intact target system state. Depending on the project, there can be WPs taking multiple inputs or generating multiple outputs. Even a uni-input/output interaction amongst the WPs however, is sufficient to convey this paper’s message.
@@ 2.3.3. Example: A structured multi-stakeholder research project
A representative case is the structure adopted in the ZORRO project, which focuses on data- and knowledge-driven diagnostics for industrial CPS (NWO:2023:ZORRO, 4). ZORRO is a typical research project focusing on algorithmic and software-based solutions. It involves academic, industrial, and applied research partners. The project is organised into seven work packages with the involvement of different users and stakeholders per work package [^1] https://zorro-project.nl/work-packages/:
@@ 4. Requirements elaboration framework
5. Discussion & example projects
@@ 5.1. The case of ZORRO
We have introduced the structure for the ZORRO project in Section 2.3.3, including its WP dependency diagram (Figure 1). With this project, there is no separate WP planning included and all available information is provided within the project proposal.
@@
More granular descriptions are provided in the WP planning, which is often appended to the project proposal itself. Specifically, the planning for WP 5 (in charge of demonstrators) reiterates the target TRL scale as 4–5. The dissemination demonstrator is described as a “simulated production environment”, covering all WPs, implying integration and thus TRL level 5. The industrial demonstrator is to apply partner use-cases, i.e., industrial partner artefact utilisation, CPP for WPs 1–4 and TFS for WP 4. Full coverage of modules using industrial artefacts points to TRL level 6. Use-case associations within the ZORRO project are given in Table 4.
@@ 5.2. The case of PrimaVera
As a second example and a project in final stages, we consider the PrimaVera project, which addresses predictive maintenance across complex systems using data-driven and cross-level optimisation methods (NWO:2019:PrimaVera, 5). It is organised into 6 technical work packages plus 2 additional packages focused on dissemination and management [^2] https://primavera-project.com/work-packages/:
@@
The WP dependency diagram for PrimaVera is given in Figure 5.
@@
Similarly to ZORRO, there is no separate WP planning included. The proposal document describes “three closely interlinked large-scale field-lab demonstrators”. The term “interlinked” indicates that even the different demonstrators are expected to implement some form of integration, which is not clarified. There are no indications of per-WP demonstrators, nor any TRL levels for any type of demonstration.
Demonstrator 1 is a health assessment and prognostics tool, which will implement the data collection, diagnostics and prognostic methods into a practical software tool. Demonstrator 2 is a planning and maintenance tool, which will implement the maintenance and supply chain optimisation into a user-friendly software tool. Demonstrator 3, the most comprehensive one, is a digital twin to support ship maintenance. The digital twin is a computer model representing the physical components and functions of (part of) a real ship, as well as its degradation behaviour. Though no TRL level is indicated, what we immediately notice from the jargon is the affinity towards tooling and targeted industrial maturity. The concept of a digital twin closely matches the high-fidelity simulated environment and pointing to TRL 7, a NASA designation (Kimmel:2020:TRAB, 6).
Further descriptions under WP 6, responsible for demonstrators, provides targeted WP coverage per each demonstrator. These coverages, alongside use-case associations within the PrimaVera project are given in Table 5. Not every single industrial partner is listed in the table. There are partners who are not clearly associated with any WP, as well as partners who do not contribute with code/data.
@@ 6. Related work
The literature often mentions terms such as “design demonstrator”, “boundary object” (Vilsmaier:2015:CMLS, 7, 8), or “intermediary objects” (Vinck:2003:EEED, 9, 10). This is partly inspired by the industry practice of new technology development and we see a “market” or “commercialisation” stage at the end. None of these term align well with our understanding from a demonstrator, as the final disseminating outcome of a research project. Most literature consider such terms as vessels of transferring artefacts between practitioners, e.g., designers and scientists.
Moultrie (Moultrie:2015:UCRD, 11) recites that in commercial setups, early involvement of industrial design has proven advantageous, leading to technologies with better applications (Gemser:2001:HIID, 12, 13). While (Moultrie:2015:UCRD, 11) relies primarily on case studies to showcase the categorisation, which is a repeating trend in other literature (Gopinath:2018:DSRI, 14, 15), the analysis itself comes across as ad hoc. A plausible reason has to be the broadness of classification. That is one motivation for us to keep the domain of applicable research projects fairly limited for the time being.

### v2

@@ Abstract.
One of the challenges apparent in the organisation of research projects is the uncertainties around the subject of demonstrators. A precise and detailed elicitation of the coverage for project demonstrators is often an afterthought and not sufficiently detailed during proposal writing. This practice leads to continuous confusion and a mismatch between targeted and achievable demonstration of results, hindering progress. The reliance on the Technology Readiness Level (TRL) scale as a loose descriptor does not help either. We propose a demonstrator requirements elaboration framework aiming to evaluate the feasibility of targeted demonstrations, making realistic adjustments, and assist in describing requirements. In doing so, we define 5 hierarchical levels of demonstration, clearly connected to expectations, e.g., work package interaction, and also connected to the project’s industrial use-cases. The considered application scope in this paper is the domain of software-intensive systems and industrial cyber-physical systems. A complete validation is not accessible, as it would require application of our framework at the start of a project and observing the results at the end, taking 4–5 years. Nonetheless, we have applied it to two research projects from our portfolio, one at the early and another at the final stages, revealing its effectiveness.
@@ 1. Introduction
Research projects often include “demonstrators” as deliverables. It is often assumed that a demonstrator is inherently known to project members, with no structure or definitive description provided as part of the project proposal. Another valid reason for the absence of such a description is the initial level of uncertainty specific to research projects. After all, a research project is supposed to be high-level by definition. Questions such as: “What is there to be demonstrated?”, “How much of it can be demonstrated?”, “Given the a priori knowledge limitations resulting from unknowns in a research project, which level of demonstration is feasible?”, are familiar to most researchers. Most projects try to solve this by relying on the Technology Readiness Levels (TRLs) (Mankins, 2004), which proves to be inadequate as it does not account for a demonstrator’s coverage. We discuss this in Section 3.4.
@@
Scope of research projects
The role of software is ever-increasing in industrial Cyber-Physical Systems (CPS), to an extent that we can consider modern instances as Software-Intensive Systems (SIS) (Hölzl:2008:ESIA; Oquendo, 2016). This designation is backed by the fact that analytic solutions intended for industrial CPS, especially the likes of solutions providing capabilities such as Fault Detection and Isolation (FDI) (Hwang et al., 2010), Fault Detection and Diagnosis (FDD) (Isermann, 2005), anomaly detection/identification (Chandola et al., 2009; Ibidunmoye et al., 2015), and predictive maintenance, are fully based on software algorithms and data. In this paper, we focus on research projects that consider use-cases from the domain of industrial CPS, with and expected outcome in the context of SIS. Further evaluations with more diverse projects could lead to adjustments, making our method suitable for a broader scope.
@@ 2. Background
2.1. TRL scale
The Technology Readiness Level (TRL) scale in its original form, was first formalised by NASA (Mankins, 2004). The scale is intended as a framework for systematic assessment of technology maturity during its development. Consisting of 9 levels from basic research (TRL 1) to fully operational (TRL 9), it was primarily focused on hardware for aerospace technologies. These levels are defined as (taken from (Mankins, 2004)):
@@ 2.3.1. Stakeholders and building blocks
  • Industrial partners usually provide real-world use-cases, validation platforms, or deployment opportunities.
@@
Demonstrator-related misalignments typically map to misunderstood, underspecified, or unspecified requirements. Especially in software-intensive projects, the absence of clearly defined interfaces, data formats, or operational contexts between WPs leads to friction during integration phases. Therefore, Requirements Engineering (RE) practices in research contexts should explicitly address targeted demonstrator scope and maturity.
@@ 2.3.2. WP dependencies
Considering the four major dependencies between software components, i.e., data dependency, control dependency, functional dependency, and temporal dependency, the most relevant dependencies for WPs can be considered as data and temporal. While we cannot dismiss control and functional dependencies, especially given the possibility of per WP software converging into a unified platform, data and temporal dependencies stand out. Incorporation of control or functional dependencies is less common as these do not align with having separate WPs. Most WPs produce output for or consume input from other WPs, which also dictates order. Order of processing can be present without any exchange as well, e.g., an analysis requiring an intact target system state. Depending on the project, there can be WPs taking multiple inputs or generating multiple outputs. Even a uni-input/output interaction amongst the WPs however, is sufficient to convey this paper’s message.
@@ 2.3.3. Example: A structured multi-stakeholder research project
A representative case is the structure adopted in the ZORRO project, which focuses on data- and knowledge-driven diagnostics for industrial CPS (NWO – Nederlandse Organisatie voor Wetenschappelijk Onderzoek, 2023). ZORRO is a typical research project focusing on algorithmic and software-based solutions. It involves academic, industrial, and applied research partners. The project is organised into seven work packages with the involvement of different users and stakeholders per work package [^1] https://zorro-project.nl/work-packages/:
@@
2.4. Targeted audience of a demonstrator
Considering the aforementioned diversity of stakeholders involved in collaborative research projects, demonstrators rarely serve a single, homogeneous audience. Each stakeholder will evaluate a given demonstrator from multiple perspectives, i.e., maturity, scope, and usefulness. As a result, the notion of a successful demonstrator is inherently stakeholder-dependent.
For instance, industrial partners prioritise clear applicability to real-world or production setups, which translates to an emphasis on robustness, scalability, system integration, and compliance with operational constraints. In contrast, academic stakeholders may focus on conceptual novelty, methodological soundness, or reusability of the demonstrator, ideally to incorporate it in future research. Funding and oversight bodies on the other hand, tend to adopt a more balanced view, searching for both scientific contribution and practical applicability.
Given the rather differing expectations, a single demonstrator can seldom satisfy all involved parties. A demonstrator optimised for either operational realism and industrial validation, or experimental freedom and exploratory research, will sacrifice opposing aspects. As such, research projects ideally should produce multiple demonstrators, or variations of a single demonstrator with explicit scopes.
Therefore, clarifying the intended audience and by extension the purpose of a demonstrator early on, is essential in research projects. Such an identification helps align expectations and results in the right set of requirements. The variability in audience expectations directly impacts how demonstrator requirements are specified and prioritised, as discussed in the following section.
2.5. Types of requirements
Software requirements can be broadly divided into functional and extra-functional categories.
Functional requirements
Functional requirements define what the system should do, specifying the behaviour, features, and interactions that enable it to perform its intended tasks. These definitions can be in varying granularities. Common examples are processing data, executing Input/Output (IO) transactions, or providing user interfaces.
Extra-functional requirements
In contrast, extra-functional requirements, a.k.a., non-functional requirements [^2] Terminology varies across the literature and can be context-dependent. Here, the term “extra-functional” is used to describe system qualities extending core functionality., define how the system should perform (Hochmuller, 1999). These requirements cover qualities such as performance metrics, reliability levels, security, maintainability, and so forth. In short, functional requirements focus on system capabilities, while extra-functional requirements define operational constraints and standards to be adhered to by the system. Software development processes intended for production-ready systems involve both types for a successful outcome.
@@ 4. Requirements elaboration framework
5. Discussion and example projects
@@ 5.1. The case of ZORRO
We have introduced the structure for the ZORRO project in Section 2.3.3, covering the projects WP descriptions and WP dependency diagram (Figure 1). With this project, there is no separate WP planning included and all available information is provided within the project proposal.
5.1.1. ZORRO WP dependency walk-through
To elaborate the potential interaction that different WPs might have, we describe the diagram from Figure 1 in a concrete scenario. The presented scenario can be considered for any industrial CPS from the project’s company use-case portfolio, e.g., an ITEC die bonder machine, or a Canon production printer:
  • WP 1’s involvement: Monitoring subsystems collect and consolidate data representing machine behaviour, covering numerous metrics, e.g., voltage, current, torque, temperature, etc., and time. The consolidation may include synthetic and semi-synthetic data generation, potentially data fusion, as well as sensor behaviour studies. For instance sensor noise simulation can be covered.
  • WP 2’s involvement: The knowledge of the system at hand is extracted and consolidated in suitable formats, e.g., a knowledge graph, from machine-specific artefacts, or expert input.
  • WP 4’s involvement: Having the domain knowledge in general and the machine’s internal operations in particular, model-based data sanitisation, augmentation, imputation, etc., is applied. Consolidated knowledge from WP 2 can be considered as an input (Figure 1), for instance describing correct data formats and hierarchical relations between different data structures, or declared associations of data segments with machine operations are to be taken into account.
  • WP 3’s involvement: Algorithms, data processing pipelines, and ML models perform Fault Detection and Isolation (FDI), consuming data passed on from WP 4. Consolidated knowledge from WP 2 is to be used in both the actual design of algorithms and from a high-level perspective, to decide on high-value faults and anomalies.
5.1.2. Considered demonstrators
@@
More granular descriptions are provided in the WP planning, which is often appended to the project proposal itself. Specifically, the planning for WP 5 (in charge of demonstrators) reiterates the target TRL scale as 4–5. The dissemination demonstrator is described as a “simulated production environment”, covering all WPs, implying integration and thus TRL level 5. The industrial demonstrator is to apply partner use-cases, i.e., industrial partner artefact utilisation, CPP for WPs 1–4 and TFS for WP 4. Full coverage of modules using industrial artefacts points to TRL level 6. Use-case associations within the ZORRO project plus a legend of the involved companies are given in Table 4.
@@ 5 CPP, TFS n/a
Legend.
   ASML: ASML Holding (https://www.asml.com/en);
   CPP: Canon Production Printing (https://cpp.canon/);
   ITEC: ITEC Equipment (https://www.itecequipment.com/);
   Philips: Philips Medical Systems (https://www.philips.nl/healthcare);
   TFS: Thermo Fisher Scientific (https://www.thermofisher.com/nl/en/home.html).
A visual diagram illustrating use-case associations at both WP level and the intended (per project proposal) industrial use-case associations within the ZORRO project, is provided in Figure 5.
Figure 5. Use-case associations at both WP and intended industrial demonstrator levels for the ZORRO project. Company use-cases with identical WP association are drawn with the same colour.
5.1.3. Applying the framework
@@ 5.2. The case of PrimaVera
As a second example and a project in its final stages, we consider the PrimaVera project, which addresses predictive maintenance across complex systems using data-driven and cross-level optimisation methods (NWO – Nederlandse Organisatie voor Wetenschappelijk Onderzoek, 2019).
5.2.1. WP descriptions
It is organised into 6 technical work packages plus 2 additional packages focused on dissemination and management [^3] https://primavera-project.com/work-packages/:
@@ 5.2.1. WP descriptions
The WP dependency diagram for PrimaVera is given in Figure 6. A concrete scenario to elaborate and crawl the diagram is given next. Note that this scenario is hypothetical and only intended to provide an understanding of the interactions.
@@
5.2.2. PrimaVera WP dependency walk-through
To elaborate the potential interaction that different WPs might have, we describe the diagram from Figure 6 in a concrete scenario. The presented scenario is specific to the naval use-case from the project and brings different WPs together around a digital twin representation of a naval vessel, as a test bench:
  • WP 1’s involvement: Data sensing algorithms, sensor network designs, data labelling strategies, and data conversion algorithms are incorporated to collect operational and maintenance data from component use-cases. Considered component data covers centrifugal pumps, bearings, vessel hull corrosion, vessel hull cracks, and impact effects.
  • WP 2’s involvement: To achieve diagnostics, several algorithms and methods are developed to detect faults, e.g., anomaly detection, or to assess the system health, e.g., impact and structural damage assessment.
  • WP 3’s involvement: Prognostic algorithms are built around three approaches: data-driven, model-based, and hybrid. Considered components cover centrifugal pumps, vessel hull corrosion, and vessel hull cracks.
  • WP 4’s involvement: Aspects such as maintenance optimisation and spare part inventory management are studied and applied to cases covering centrifugal pumps, bearings, and vessel hull corrosion.
  • WP 5’s involvement: User-friendly decision support algorithm design and organisational implementation strategies are covered and applied to centrifugal pumps, vessel hull corrosion, and vessel hull cracks.
5.2.3. Considered demonstrators
Similar to the ZORRO project, there is no separate WP planning included. The proposal document describes “three closely interlinked large-scale field-lab demonstrators”. The term “interlinked” indicates that even the different demonstrators are expected to implement some form of integration, which is not clarified. There are no indications of per-WP demonstrators, nor any TRL levels for any type of demonstration.
Demonstrator 1 is a health assessment and prognostics tool, which will implement the data collection, diagnostics and prognostic methods into a practical software tool. Demonstrator 2 is a planning and maintenance tool, which will implement the maintenance and supply chain optimisation into a user-friendly software tool. Demonstrator 3, the most comprehensive one, is a digital twin to support ship maintenance. The digital twin is a computer model representing the physical components and functions of (part of) a real ship, as well as its degradation behaviour. Though no TRL level is indicated, what we immediately notice from the jargon is the affinity towards tooling and targeted industrial maturity. The concept of a digital twin closely matches the high-fidelity simulated environment and pointing to TRL 7, a NASA designation (Kimmel et al., 2020).
Further descriptions under WP 6, responsible for demonstrators, provides targeted WP coverage per each demonstrator. These coverages, alongside use-case associations within the PrimaVera project are given in Table 5. A legend of the involved companies is included as well. Not every single industrial partner is listed in the table. There are partners who are not clearly associated with any WP, as well as partners who do not contribute with code/data.
@@ 5.2.3. Considered demonstrators
Legend.
   Alfa Laval: Alfa Laval Benelux (https://www.alfalaval.nl/nl/);
   ASML: ASML Holding (https://www.asml.com/en);
   Damen: Damen Shipyards Group (https://www.damen.com/);
   IHC: Royal IHC (https://www.royalihc.com/);
   NS: Netherlands Railways (https://www.ns.nl/en);
   RNL: Royal Netherlands Navy (https://english.defensie.nl/organisation/navy);
   RWS: Rijkswaterstaat (https://www.rijkswaterstaat.nl/en);
   WDD: Water Board de Dommel (https://www.dommel.nl/).
A visual diagram illustrating use-case associations at both WP level and the intended (per project proposal) industrial use-case associations within the PrimaVera project, is provided in Figure 7.
Figure 7. Use-case associations at both WP and intended industrial demonstrator levels for the PrimaVera project. Company use-cases with identical WP association are drawn with the same colour.
@@ 6. Related work
The literature often mentions terms such as “design demonstrator”, “boundary object” (Vilsmaier et al., 2015; Feldhoff et al., 2019), or “intermediary objects” (Vinck, 2003, 2012). This is partly inspired by the industry practice of new technology development and we see a “market” or “commercialisation” stage at the end. None of these term align well with our understanding from a demonstrator, as the final disseminating outcome of a research project. Most literature consider such terms as vessels of transferring artefacts between practitioners, e.g., designers and scientists.
Moultrie (Moultrie, 2015) recites that in commercial setups, early involvement of industrial design has proven advantageous, leading to technologies with better applications (Gemser and Leenders, 2001; Hertenstein et al., 2001). While (Moultrie, 2015) relies primarily on case studies to showcase the categorisation, which is a repeating trend in other literature (Gopinath et al., 2018; Bobbe et al., 2023), the analysis itself comes across as ad hoc. A plausible reason has to be the broadness of classification. That is one motivation for us to keep the domain of applicable research projects fairly limited for the time being.